OpenStudy (anonymous):

Find the indefinite intergral: sqrtx+1/(2sqrtx)

OpenStudy (anonymous):

This? \[\int\limits \frac{\sqrt{x+1}}{2\sqrt{x}}dx\]

OpenStudy (anonymous):

No... interal of sqaure root of x + (1/(2sqrtx)) \[\int\limits\]

OpenStudy (anonymous):

\[\int\limits \sqrt{x} + \frac{1}{2\sqrt{x}}dx \space \space ?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, well that's a lot easier then. You can break it up and rewrite it as: \[\int\limits x^{1/2}dx+\frac{1}{2}\int\limits x^{-1/2}dx\]

OpenStudy (anonymous):

Then use power rule.

OpenStudy (anonymous):

Ok thanks so much, could you figure out the final answer to check it with mine

OpenStudy (anonymous):

i got: \[1/3 \sqrt{x} (2x+3)+C\]

OpenStudy (anonymous):

Sure, just a sec.

OpenStudy (anonymous):

No. How did you get that?

OpenStudy (anonymous):

\[\int\limits x^{1/2}dx = \frac{2}{3}x^{3/2}+C.\]